# Expanding Horizons in Short Text Analysis: Integrating LLMs and VAEs for Enhanced Topic Modeling

### v2

@@ 1 Introduction
In recent years, the field of natural language processing has witnessed remarkable advancements recently, largely attributable to the proliferation of Large Language Models (LLMs). These LLMs, characterized by their extensive scale training on vast corpora, have demonstrated unparalleled capability of understanding human text and producing high-quality responses kaplan2020scaling; clark2022unified. Moreover, they showcase exceptional practicality and scalability across a spectrum of downstream NLP tasks, such as conversational response generation, text summarization, or content recommendation, under only few-shot or zero-shot settings wu2023survey. However, despite their potential, a major limitation of these LLMs is their struggle to accurately capture the dynamic nature of information updates. Since LLMs are pre-trained on static corpora of data they face difficulties when dealing with the latest or less commonly known information that falls outside their initial training scope. This limitation leads to generating responses that contain outdated or incorrect information, or in some cases, entirely fictitious content— a problem often termed as "hallucination" lee2018hallucinations.
In response to this pressing challenge, the concept of Retrieval-Augmented Generation (RAG) RAG has emerged as a promising technique to enhance the functionality and reliability of LLMs. Retrieval augmentation addresses the shortcomings of original LLMs by incorporating external knowledge (i.e., non-parametric knowledge) sources through some information retrieval system and leveraging the in-context learning ability of LLM readers to generate more contextually relevant and accurate responses. For instance, in a conversational setting where a user interacts with an LLM like ChatGPT gpt to inquire about the latest news events. ChatGPT, being reliant on static pre-training knowledge, may struggle to offer real-time updates in response to the user query. RAG mitigates this information gap by retrieving pertinent news articles or data related to the query from external databases. By augmenting the original user question with retrieved knowledge, the final LLM reader is more likely to generate a contextually relevant answer.
While previous research has underscored the efficacy of retrieval augmentation in mitigating hallucination, it introduces its own set of challenges. One such critical challenge in the RAG systems is the existence of a gap between the information retrievable using the given retriever using original user input queries and the information required for generating optimal responses ragsurvey. Consider a scenario where an external document collection comprises three distinct passages labelled Passage $A$, $B$, and $C$, each containing unique knowledge components denoted as knowledge $x$, $y$, and $z$, respectively. Notably, all three passages include keywords related to Knowledge $z$, which the user seeks to obtain. However, if the user enters a poorly formulated query, there exists a possibility that Passage $A$ or Passage $B$ may be retrieved instead of the ideal source, Passage $C$, thereby limiting the LLM reader’s capacity to derive an optimal response. This disparity is referred to as the pre-retrieval gap in RAG and poses a fundamental obstacle to achieving optimal text generation outcomes.
In the quest to bridge this pre-retrieval information gap ragsurvey within the existing RAG systems, the Rewrite-Retrieve-Read (RRR) framework rrr introduces a pivotal paradigm shift by incorporating the concept of query rewriting from the realm of traditional information retrieval into the RAG domain. RRR proposes the integration of a specialized Query Rewriter module into the native RAG system to align user queries with both the retriever and the LLM reader, thereby optimizing the retrieval process and narrowing the pre-retrieval gap between user queries and the information sought by the LLM reader, i.e. the LLM that generates the final responses. This query optimization technique has also seen widespread adoption in more recent RAG systems stepback; ragsurvey. However, while RRR demonstrates notable advancements, only the original query is passed to the rewrite module, which focuses more on rephrasing queries or expanding the search scope. This framework falls short in tailoring optimization strategies for queries based on the specific needs of the LLM reader, overlooking the need for retrieving the most relevant knowledge essential for accurate generation.
@@
To this end, we propose Extract-Refine-Retrieve-Read (ERRR), a simple but effective framework designed for retrieval augmentation systems. The ERRR framework is crafted to bridge the pre-retrieval information gap through tailored query optimization and aims to resolve the inherent limitations of RRR by enabling retrieval based on the specific information needs of the LLM reader. Specifically, it initiates by extracting parametric knowledge from LLMs and employs a specialized query optimizer that refines user queries. This refinement either complements or validates the extracted parametric knowledge, ensuring that only essential information is retrieved for generating accurate responses, and minimizing the retrieval of extraneous information that could degrade the quality of the output.
Additionally, given that many advanced LLMs, such as ChatGPT gpt, function as black-box systems accessible only through inference APIs, our study introduces a trainable ERRR scheme to enhance customizability for different types of queries. This scheme incorporates a smaller, tunable language model as the query optimizer, which not only helps to lower the expenses associated with query optimization for parametric knowledge refinement but also allows for more adaptable customization tailored to each knowledge source.
Our methods undergo thorough evaluation on multiple question-answering (QA) datasets (HotpotQA hotpotqa, AmbigNQ ambigqa, PopQA popqa). All frameworks and baselines evaluated are implemented using GPT-3.5-Turbo gpt3 or T5-Large t5 as the query optimizer, with GPT-3.5-Turbo serving as both the final reader and the source of parametric knowledge. Evaluation is conducted on the Brave Web Search Engine  [^1] https://brave.com/search/api/ and a local dense retriever, Dense Passage Retrieval (DPR) DPR, alongside a static open-source Wikipedia corpus. Our evaluations demonstrate that ERRR consistently boosts the performance of retrieval-augmented LLMs across all tested QA datasets and retrieval systems when compared to the RRR framework. Furthermore, the outcomes highlight the adaptability and versatility of ERRR, showcasing its effectiveness across diverse settings and data sources.
In summary, our key contributions are as follows: (i) We propose Extract-Refine-Retrieve-Read (ERRR), a retrieval augmentation framework designed to boost the effectiveness of RAG systems by optimizing queries to meet the specific knowledge needs of LLMs. (ii) We reveal that ERRR has a high degree of adaptability across diverse settings and data sources, as well as across various retrieval systems. (iii) We introduce a trainable scheme for ERRR, which not only achieves higher performance but also effectively reduces computational costs.
@@ 2.1 Retrieval-Augmented Generation
The integration of retrieval modules to access relevant contextual knowledge has played a crucial role in enhancing Large Language Models (LLMs) in recent years. Initially designed for early sequence-to-sequence models, the Retrieval-Augmented Generation (RAG) framework proposed by Piktus et al. RAG has gained substantial traction in the era of LLM. This approach has diversified into a broad array of methods, with ongoing efforts aimed at further augmenting its capabilities. Earlier exploration primarily focused on improving key components, such as upgrading to more powerful pre-trained language models like BERT bert as readers or employing advanced dense retrievers for retrieval tasks DPR. These retrievers encode documents and inputs into dense vectors, facilitating retrieval based on the similarity between the input and retrieved passages.
Recent studies have shifted focus beyond merely enhancing the retriever or reader components, emphasizing the refinement of pre-retrieval and post-retrieval processes. To address the pre-retrieval gap—the disparity between the information retrievable from original queries and the knowledge required for optimal responses—GenRead genread replaces the retrieval module with a knowledgeable LLM, thereby narrowing the gap between the user query and retrieval process. It prompts the LLM to generate contextual information for the query, using these generated documents as retrieval results to formulate the final answer. Self-ask selfask proposes an iterative approach using chain-of-thought prompting to generate self-posed questions that refine the response. For the post-retrieval gap—the challenge of creating optimal responses from given information—strategies include document re-ranking or summarization. For instance, PRCA prca trains a contextual adaptor module to summarize retrieved documents with a black-box LLM reader.
Several studies have also proposed significant modifications to the original RAG pipeline, introducing complex systems that include both pre-retrieval and post-retrieval modules ragfusion, and adapting the pipeline into iterative or recursive frameworks react; selfrag. While these advanced systems demonstrate notable performance enhancements, they incur substantial costs and typically require multiple interactions with LLM. In contrast, our work focuses on refining the single-turn RAG framework, introducing a flexible and trainable module adaptable to existing systems.
@@ 2.2 Query Optimization for Retrieval Augmentation
Recent research highlights a significant discrepancy between input queries and LLM readers for RAG systems, especially under the current trend of using off-the-shelf web search tools or black-box LLMs that are difficult to customize rrr. Typically, these input queries often originate directly from users or specific datasets, which could be either poorly formulated or adhere to a static query format. To overcome these challenges, an effective approach is to optimize the query in the pre-retrieval phase, thereby improving the quality of retrieved information and response generation. The Rewrite-Retrieve-Read (RRR) framework, for instance, trains a query rewriting module using an LLM to better align retrieval queries with LLM readers rrr that generate the final response, as illustrated in Figure 1. Additionally, RRR introduces a trainable scheme that employs reinforcement learning with Proximal Policy Optimization to fine-tune a small open-source model based on feedback from the LLM reader, achieving improved results. HyDE addresses the demand for accurate information retrieval by creating hypothetical documents and encoding them through unsupervised contrastive learning for efficient retrieval operations hyde. Furthermore, Step-Back Prompting stepback converts original queries into high-level abstract questions, aiding LLMs in generating better responses for complex queries requiring abstract thinking.
@@ 3.1 Pre-retrieval Information Gap
[…]
@@
Previous work like RRR rrr has demonstrated the effectiveness of such query optimization functions, albeit without considering the influence of $\theta$. To this end, ERRR introduces a more tailored query optimization function $f^{\prime}$ that utilizes the parametric knowledge $\theta$ to perform the query optimization and retrieve external knowledge that refines $\theta$ and better aligns with its needs. This can be formulated as:
[…]
@@ 3.2 Extract-Refine-Retrieve-Read
Parametric Knowledge Extraction  Previous studies such as GenRead genread and HyDE hyde demonstrate that LLMs may possess substantial parametric knowledge capable of addressing user inquiries, particularly on popular topics. Inspired by the prompting methods outlined in GenRead, our approach involves a direct strategy where we prompt the LLM reader to produce a pseudo-contextual document containing all the background information. We consider these pseudo-contextual documents as a representation of the LLM’s abstracted parametric knowledge. Although these documents may contain inaccuracies, they provide essential contextual information related to the original queries.
@@ 3.3 Trainable Scheme
{tabularx}
[…]
@@ 4.1 Datasets and Metrics
ERRR is assessed on three open-domain question-answering (QA) datasets: AmbigQA ambigqa, PopQA popqa, and HotpotQA hotpotqa. Each dataset serves to test different capabilities of the ERRR framework. (i) The AmbigNQ dataset is the disambiguated variant of the Natural Questions (NQ) dataset, where ambiguous questions from NQ are refined into specific queries with minimal constraints. Consistent with procedures used in RRR, we evaluated ERRR using the first 1000 samples of the test set. (ii) PopQA features simpler questions that focus on less popular knowledge topics compared to other QA tasks. Due to the high similarity in sample distributions, we assessed only the first 997 samples of the test set. (iii) The HotPotQA dataset contains complex questions that require multi-hop reasoning. We conducted evaluations across the entire test set. Following the metric usage for three datasets, our method is evaluated by exact match score $EM$ and $F_{1}$ score.
@@ 4.2 Baselines and Proposed Frameworks
We evaluated 7 baselines and proposed frameworks, as detailed below: (i) Direct: Directly calling GPT-3.5-Turbo to answer questions. (ii) RAG: The classic Retrieval-Augmented Generation framework RAG. The original user queries are used for retrieval and fed directly to the LLM reader to generate output. (iii) ReAct: A modified RAG framework that intertwines the reasoning and acting capabilities of LLMs to create a more cohesive and effective approach react. This framework can iteratively perform reasoning prompts and actions, such as information retrieval, serving as our comparison baseline. (iv) Frozen RRR: Rewrite-Retrieve-Read framework rrr with a frozen configuration. It employs GPT-3.5-Turbo to rewrite the query and retrieve relevant documents based on these rewritten queries. Then the original query and retrieved documents are used for reading. This serves as our baseline for comparison. (v) Trainable RRR: Trainable rewrite-retrieve-read framework, initiating with a supervised fine-tuned T5-large model. It then applies reinforcement learning to better align the retriever and rewriter using Proximal Policy Optimization (PPO). This serves as our baseline for comparison. (vi) Frozen ERRR: Extract-Refine-Retrieve-Read framework with a frozen configuration, as described in Section 3.2. (vii) Trainable ERRR: Trainable Extract-Refine-Retrieve-Read framework, as described in Section 3.3.

### v3

# Query Optimization for Parametric Knowledge Refinement in Retrieval-Augmented Large Language Models

@@ Abstract
Query Optimization for Parametric Knowledge Refinement in Retrieval-Augmented Large Language Models
Youan Congthanks: Equal contribution.   Pritom Saha Akash11footnotemark: 1   Cheng Wang   Kevin Chen-Chuan Chang University of Illinois Urbana-Champaign, USA {youanc2, pakash2, chengw4, kcchang}@illinois.edu
@@ 1 Introduction
The field of natural language processing (NLP) has witnessed transformative advancements in recent years, largely driven by the advent of Large Language Models (LLMs). These models, trained on vast corpora, have demonstrated exceptional capabilities in understanding human text and generating high-quality responses Kaplan et al. (2020); Clark et al. (2022). They have also proven practical and scalable for various downstream NLP tasks, such as conversational response generation, text summarization, and content recommendation, even in few-shot or zero-shot settings Wu et al. (2023). Despite their strengths, a key limitation of LLMs lies in their reliance on static training data, which makes them struggle with dynamic or less commonly known information outside their initial training scope. This limitation often leads to outdated, inaccurate, or entirely fabricated responses—a phenomenon commonly referred to as “hallucination” Lee et al. (2018).
To address this issue, Retrieval-Augmented Generation (RAG) Lewis et al. (2020) has emerged as a promising approach to enhance the functionality and reliability of LLMs. By integrating external knowledge sources through retrieval systems, RAG enables LLMs to augment user queries with relevant, up-to-date information. This augmentation allows LLMs to generate more contextually accurate and relevant responses. For instance, in a conversational setting where a user queries an LLM like ChatGPT Ouyang et al. (2022) for the latest news, RAG retrieves pertinent articles to supplement the static pre-trained knowledge of the model, thereby mitigating the information gap.
While retrieval augmentation has proven effective in mitigating hallucinations, it introduces its own set of challenges. A prominent issue in Retrieval-Augmented Generation (RAG) systems is the pre-retrieval gap—a mismatch between the information retrieved using the original user query and the specific knowledge required to generate optimal responses Gao et al. (2024). For instance, consider a document collection containing three passages, labeled Passage $A$, $B$, and $C$, each containing unique knowledge components $x$, $y$, and $z$, respectively. Although all three passages include keywords associated with Knowledge zz—the user’s intended target—a poorly formulated query may lead to retrieving Passage $A$ or $B$ instead of the ideal Passage $C$. This misalignment restricts the LLM reader’s ability to generate accurate responses, making the pre-retrieval gap a critical barrier to achieving optimal text generation in RAG systems.
To bridge the pre-retrieval gap, the Rewrite-Retrieve-Read (RRR) framework Ma et al. (2023) introduced query rewriting as a mechanism to optimize user queries and improve their alignment with retrieval systems. However, RRR and similar methods Zheng et al. (2024); Gao et al. (2024) primarily focus on rephrasing or broadening queries, which helps expand the search scope but fails to address the specific knowledge requirements of the LLM reader. Additionally, recent self-prompting techniques Li et al. (2022); Wang et al. (2023) have explored using chain-of-thought (CoT) prompts and pseudo-QA pairs to enhance LLM reasoning capabilities by eliciting internal parametric knowledge. While these approaches effectively improve the internal reasoning and explanation capabilities of LLMs for tasks like multi-hop reasoning and open-domain QA, they lack mechanisms for aligning external retrieval queries with the LLM’s knowledge gaps, making them insufficient for resolving the pre-retrieval gap in Retrieval-Augmented Generation (RAG) systems.
@@
To this end, we propose Extract-Refine-Retrieve-Read (ERRR), a simple but effective framework designed for retrieval augmentation systems. The ERRR framework is crafted to bridge the pre-retrieval information gap through tailored query optimization and aims to resolve the inherent limitations of RRR by enabling retrieval based on the specific information needs of the LLM reader. Specifically, it initiates by extracting parametric knowledge from LLMs and employs a specialized query optimizer which refines user queries. This refinement either complements or validates the extracted parametric knowledge, ensuring that only essential information is retrieved for generating accurate responses, and minimizing the retrieval of extraneous information that could degrade the quality of the output.
In addition to its innovative query optimization process, ERRR introduces a trainable scheme to enhance efficiency and adaptability. Recognizing the constraints posed by black-box systems like ChatGPT Ouyang et al. (2022), which are accessible only through inference APIs, ERRR incorporates a smaller, tunable language model as the query optimizer. This trainable component reduces computational costs while offering greater flexibility to customize the retrieval process for diverse queries and knowledge sources. By combining precision in addressing pre-retrieval gaps with cost-effective adaptability, ERRR provides a robust solution for improving retrieval augmentation in LLM-driven systems.
We evaluate ERRR on multiple question-answering (QA) datasets, including HotpotQA Yang et al. (2018), AmbigNQ Min et al. (2020), and PopQA Mallen et al. (2022), using both web-based (e.g., Brave Search Engine) and local retrieval systems (e.g., Dense Passage Retrieval Karpukhin et al. (2020)). Across all tested datasets and retrieval configurations, ERRR consistently outperforms baseline frameworks, such as RRR, in terms of retrieval accuracy and response quality. These results highlight ERRR’s versatility and effectiveness in diverse settings.
In summary, our key contributions are as follows: (i) We propose Extract-Refine-Retrieve-Read (ERRR), a novel framework that optimizes queries to bridge the pre-retrieval gap and enhance RAG systems. (ii) We demonstrate ERRR’s adaptability across different datasets, retrieval systems, and settings, establishing its robustness and versatility. (iii) We introduce a trainable ERRR scheme that reduces computational costs while maintaining high performance, making it suitable for real-world applications.
@@ 2.1 Retrieval-Augmented Generation
The integration of retrieval modules to access relevant contextual knowledge has played a crucial role in enhancing Large Language Models (LLMs) in recent years. Initially designed for early sequence-to-sequence models, the Retrieval-Augmented Generation (RAG) framework proposed by Piktus et al. Lewis et al. (2020) has gained substantial traction in the era of LLM. This approach has diversified into a broad array of methods, with ongoing efforts aimed at further augmenting its capabilities. Earlier exploration primarily focused on improving key components, such as upgrading to more powerful pre-trained language models like BERT Devlin et al. (2019) as readers or employing advanced dense retrievers for retrieval tasks Karpukhin et al. (2020). These retrievers encode documents and inputs into dense vectors, facilitating retrieval based on the similarity between the input and retrieved passages.
Recent studies have shifted focus beyond merely enhancing the retriever or reader components, emphasizing the refinement of pre-retrieval and post-retrieval processes. To address the pre-retrieval gap—the disparity between the information retrievable from original queries and the knowledge required for optimal responses—GenRead Yu et al. (2023) replaces the retrieval module with a knowledgeable LLM, thereby narrowing the gap between the user query and retrieval process. It prompts the LLM to generate contextual information for the query, using these generated documents as retrieval results to formulate the final answer. Self-ask Press et al. (2023) proposes an iterative approach using chain-of-thought prompting to generate self-posed questions that refine the response. For the post-retrieval gap—the challenge of creating optimal responses from given information—strategies include document re-ranking or summarization. For instance, PRCA Yang et al. (2023) trains a contextual adaptor module to summarize retrieved documents with a black-box LLM reader.
Several studies have also proposed significant modifications to the original RAG pipeline, introducing complex systems that include both pre-retrieval and post-retrieval modules Rackauckas (2024), and adapting the pipeline into iterative or recursive frameworks Yao et al. (2022); Asai et al. (2023). While these advanced systems demonstrate notable performance enhancements, they incur substantial costs and typically require multiple interactions with LLM. In contrast, our work focuses on refining the single-turn RAG framework, introducing a flexible and trainable module adaptable to existing systems.
@@ 2.2 Query Optimization for Retrieval Augmentation
Recent research highlights a significant discrepancy between input queries and LLM readers for RAG systems, especially under the current trend of using off-the-shelf web search tools or black-box LLMs that are difficult to customize Ma et al. (2023). Typically, these input queries often originate directly from users or specific datasets, which could be either poorly formulated or adhere to a static query format. To overcome these challenges, an effective approach is to optimize the query in the pre-retrieval phase, thereby improving the quality of retrieved information and response generation. The Rewrite-Retrieve-Read (RRR) framework, for instance, trains a query rewriting module using an LLM to better align retrieval queries with LLM readers Ma et al. (2023) that generate the final response, as illustrated in Figure 1. Additionally, RRR introduces a trainable scheme that employs reinforcement learning with Proximal Policy Optimization to fine-tune a small open-source model based on feedback from the LLM reader, achieving improved results. HyDE addresses the demand for accurate information retrieval by creating hypothetical documents and encoding them through unsupervised contrastive learning for efficient retrieval operations Gao et al. (2023). Furthermore, Step-Back Prompting Zheng et al. (2024) converts original queries into high-level abstract questions, aiding LLMs in generating better responses for complex queries requiring abstract thinking.
@@ 3.1 Pre-retrieval Information Gap
$\text{LLM}(D,q~|~\theta)=a$ (1)
where LLM denotes an LLM reader and $\theta$ denotes the parametric knowledge of the LLM.
@@
Previous work like RRR Ma et al. (2023) has demonstrated the effectiveness of such query optimization functions, albeit without considering the influence of $\theta$. To this end, ERRR introduces a more tailored query optimization function $f^{\prime}$ that utilizes the parametric knowledge $\theta$ to perform the query optimization and retrieve external knowledge that refines $\theta$ and better aligns with its needs. This can be formulated as:
$\text{LLM}(R(f^{\prime}(C,q)),q~|~\theta)$ (4)
@@ 3.2 Extract-Refine-Retrieve-Read
Parametric Knowledge Extraction  Previous studies such as GenRead Yu et al. (2023) and HyDE Gao et al. (2023) demonstrate that LLMs may possess substantial parametric knowledge capable of addressing user inquiries, particularly on popular topics. Inspired by the prompting methods outlined in GenRead, our approach involves a direct strategy where we prompt the LLM reader to produce a pseudo-contextual document containing all the background information. We consider these pseudo-contextual documents as a representation of the LLM’s abstracted parametric knowledge. Although these documents may contain inaccuracies, they provide essential contextual information related to the original queries.
@@ 3.3 Trainable Scheme
Direct Prompt
Answer the question in the following format, end the answer with ’**’. {demonstration} Question: {xx} Answer:
Reader Prompt for Retrieval Augmentation Generation
Answer the question in the following format, end the answer with ’**’. {demonstration} Question: {docdoc} {xx} Answer:
Prompt for RRR Query Rewriter
Think step by step to answer this question, and provide search engine queries for knowledge that you need. Split the queries with ’;’ and end the queries with ’**’. {demonstration} Question: {xx} Answer:
Prompt for Parametric Knowledge Extraction
Generate a background document from web to answer the given question. {xx}
Prompt for ERRR Query Optimizer
Address the following questions based on the contexts provided. Identify any missing information or areas requiring validation, especially if time-sensitive data is involved. Then, formulate several specific search engine queries to acquire or validate the necessary knowledge. Split the queries with ’;’ and end the queries with ’**’. {demonstration} Context: {Parametric Knowledge} Question: {xx} Queries:
@@ 4.1 Datasets and Metrics
ERRR is assessed on three open-domain question-answering (QA) datasets: AmbigQA Min et al. (2020), PopQA Mallen et al. (2022), and HotpotQA Yang et al. (2018). Each dataset serves to test different capabilities of the ERRR framework. (i) The AmbigNQ dataset is the disambiguated variant of the Natural Questions (NQ) dataset, where ambiguous questions from NQ are refined into specific queries with minimal constraints. Consistent with procedures used in RRR, we evaluated ERRR using the first 1000 samples of the test set. (ii) PopQA features simpler questions that focus on less popular knowledge topics compared to other QA tasks. Due to the high similarity in sample distributions, we assessed only the first 997 samples of the test set. (iii) The HotPotQA dataset contains complex questions that require multi-hop reasoning. We conducted evaluations across the entire test set. Following the metric usage for three datasets, our method is evaluated by exact match score $EM$ and $F_{1}$ score.
@@ 4.2 Baselines and Proposed Frameworks
We evaluated 7 baselines and proposed frameworks, as detailed below: (i) Direct: Directly calling GPT-3.5-Turbo to answer questions. (ii) RAG: The classic Retrieval-Augmented Generation framework Lewis et al. (2020). The original user queries are used for retrieval and fed directly to the LLM reader to generate output. (iii) ReAct: A modified RAG framework that intertwines the reasoning and acting capabilities of LLMs to create a more cohesive and effective approach Yao et al. (2022). This framework can iteratively perform reasoning prompts and actions, such as information retrieval, serving as our comparison baseline. (iv) Frozen RRR: Rewrite-Retrieve-Read framework Ma et al. (2023) with a frozen configuration. It employs GPT-3.5-Turbo to rewrite the query and retrieve relevant documents based on these rewritten queries. Then the original query and retrieved documents are used for reading. This serves as our baseline for comparison. (v) Trainable RRR: Trainable rewrite-retrieve-read framework, initiating with a supervised fine-tuned T5-large model. It then applies reinforcement learning to better align the retriever and rewriter using Proximal Policy Optimization (PPO). This serves as our baseline for comparison. (vi) Frozen ERRR: Extract-Refine-Retrieve-Read framework with a frozen configuration, as described in Section 3.2. (vii) Trainable ERRR: Trainable Extract-Refine-Retrieve-Read framework, as described in Section 3.3.